The length of a rectangle is four times its width.
If the perimeter of the rectangle is 90ft , find its area.

Answer by TimothyLamb(4379)

L = 4W
P = 2L + 2W = 90
---
2L + 2W = 90
2(4W) + 2W = 90
8W + 2W = 90
10W = 90
---
W = 9
L = 36
---
A = 324
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php